OpenStudy (anonymous):

Use an appropriate Half-Angle Formula to find the exact value of the expression. sin(7π/8)????

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \tt π=180° }\) \(\Huge\color{blue}{ \tt 7π/8=7(180)/8=?}\)

OpenStudy (solomonzelman):

you tell me

OpenStudy (solomonzelman):

i forgot the formula for half angle for sin

OpenStudy (anonymous):

\[\sin(\theta/2)= \pm \frac {\sqrt{1-\cos(\theta)}} {2}\] \]

OpenStudy (solomonzelman):

\[\sin(\theta/2)= \pm \frac {\sqrt{1-\cos(\theta)}} {2}\]

OpenStudy (solomonzelman):

you angle is 157.5 so... \(\Huge\color{blue}{ \sf sin \frac{315}{2} = ±\frac{ \sqrt{1-cos(315)} }{2} }\) for cos315, use cos(315)=cos(360-45)

OpenStudy (solomonzelman):

cos(A-B)=cos A cos B + sin A sin B

OpenStudy (anonymous):

and the 315 came from where ?

OpenStudy (solomonzelman):

because you are using half angle formula, and 157.5=315/2

OpenStudy (solomonzelman):

cos(A-B)=cos A cos B + sin A sin B cos(315) = cos(360-45) = cos 360 cos 45 + sin 360 sin 45 = 1 * (sqrt 2)/2 + 0 * (sqrt 2)/2 = (sqrt 2)/2

OpenStudy (anonymous):

\[1-\sqrt{2}/2\]/2

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf Sin\frac{315}{2}= \frac{\sqrt{1- \frac{ \sqrt{2} }{2} } }{2} }\) \(\Huge\color{blue}{ \sf Sin\frac{315}{2}= \frac{\sqrt{\frac{2}{2}- \frac{ \sqrt{2} }{2} } }{2} }\) \(\Huge\color{blue}{ \sf Sin\frac{315}{2}= \frac{\sqrt{ \frac{ 2-\sqrt{2} }{2} } }{2} }\) hold....